# On the Use of WGANs for Super Resolution in Dark-Matter Simulations

### v1

@@ 3.1 Enhanced Dark-Matter Snapshot
Figure 5: The halo mass function was computed using a friends-of-friends algorithm, with a linking length of $0.2$ times the mean inter-particle spacing, to identify halos. Takeaway: The halo mass function of the enhanced snapshot closely matches that of the HR simulation, including at mass scales not resolved by the LR snapshot.
@@
The halo mass function for each snapshot was also computed using a friends-of-friends method (e.g. Davis et al., 1985) with a linking length of $0.2$ times the mean inter-particle spacing and by only considering halos with at least 100 particles. We similarly find that the enhanced data agrees with the HR data. The mass functions are shown in Fig. 5 and demonstrate that the enhanced snapshot successfully reproduces the abundance of halos across the same mass range as the HR data. In particular, the enhanced snapshot captures the higher abundance of low-mass halos, which are poorly resolved in the LR data, while maintaining accuracy for high-mass halos. These results indicate that the enhancement method not only improves the resolution of the displacement fields but also faithfully reproduces critical statistical properties of the HR simulations.
@@
Enhancing LR snapshots with a trained generator model takes about 2 seconds on a single CPU  [^1] The CPU used was Intel Core i9-10940X CPU. This is negligible compared to the wall-clock time for the LR simulation to reach redshift 00 using the same CPU (approximately 7 minutes). While the wall-clock time for the HR counter part to reach the same redshift was 1 hour, generating HR snapshots with SR is significantly faster, with greater speed-ups expected when larger enhancement factors are used.HR-comparedHR The smearing of halo shapes appears to occur on length scales below the spatial resolution of the density fields used during training. As a result, the smeared features are not resolved by the cloud-in-cells density fields used during training, meaning the neural network models may learn to ignore it. This discrepancy could pose challenges for any SR applications where analysis is sensitive to such smearing.
@@
In principle, this issue can be addressed by increasing the resolution of the density fields used during training. However, we find that the increase in resolution needed to remove this smearing effect is computationally prohibitively expensive. To test this approach, we trained two additional models using the generalized architecture for the critic model described in section 2.2.2. The generalized critic model facilitated training WGAN models using density fields with spatial dimensions $68^{3}$ and $140^{3}$, corresponding to one and two additional residual blocks in the density branch of the critic model respectively. The base number of channels for the density branch in both cases was set to $16$ while it was kept at $64$ for the main branch. A comparison of dark-matter distributions generated by these models is shown in Fig 8 in the appendix. We found that the increases in resolution to the density field that we tested were not sufficient to remove the smearing effect from generated data. Furthermore, the wallclock time needed to perform a single epoch increased by a factor of 2 and 140 respectively. Larger increases to the density resolution were not tested as the increased memory footprint exceeded the capacity of the H100 GPU device used for this work.
@@ 4 Conclusions and Outlook
In this work we trained several WGAN models to increase the particle resolution of snapshots from dark-matter-only simulations using the method outlined in Li et al. (2021). We reproduced results that confirm that these models are capable of converting LR dark-matter data into HR data with summary statistics that agree with real HR simulations. Furthermore, plots of the generated dark-matter density distribution are difficult to distinguish visually from those coming from real HR simulations. However, we found a notable limitation: HR snapshots generated by these models exhibit smeared features that clearly distinguish them from their real HR counterparts.
@@
Despite these limitations, the SR method we tested, along with similar techniques, remain a valuable tool for cosmological studies. Given the impressive demonstrations of the techniques used in Zhang et al. (2024) and Jacobus et al. (2024), it is tempting to judge these methods as cheap replacements for HR simulations. Nevertheless, it is important to recognize that these models are approximating the physics occurring on length scales that are not resolved by the LR simulation. Therefore, rather than being viewed as a replacement for HR simulations, SR methods should be regarded as simulation-informed subgrid models. As such, the value of these methods stems from their ability to help bridge the gap between LR and HR simulations when computational resources are limited.
@@
Future work could address the issue of smeared features by developing additional loss terms that account for physical properties beyond summary statistics. For example, quantifying the divergence of halo shapes from those appearing in HR data could provide a regularization term to discourage the WGAN from generating overly smoothed solutions. This approach aligns with the methods used by Helfer et al. (2024) to correct numerical relativity simulations that break conservation laws during mesh refinement. Exploring alternative generative frameworks, such as stable diffusion models, may also provide insights into whether similar limitations persist or if new architectures can overcome this issue. A detailed comparison between WGANs and stable diffusion methods in this context could help identify the strengths and weaknesses of each approach.
@@
Applications of SR techniques are significant, and we will explore some of those applications in future works. One example would be the ability to extend the resolution of cosmological N-body simulations to smaller scales opens the possibility of better tracking the assembly history of halos at a relatively small cost. In addition, to the positional information used in this work, this would require temporal (redshift) information and kinematics - however both of these extensions are tractable. Gaining insight into the assembly history of halos at high redshift would be particularly useful for understanding the formation of (rare) early cosmic structures which require both large volume statistics and high resolution (e.g. Trinca et al., 2022). These smaller halos (atomic cooling halos and molecular cooled halos) have the potential to host the first stellar objects in our Universe. Semi-analytical models, which use merger trees as their backbone, could benefit hugely from SR techniques (e.g. Spinoso et al., 2023).
@@
A second example is the use of SR techniques to capture small scale information in cosmological structure formation simulations on scales of 100s Mpc in realizations of large-scale structure on several Gpc scales. The shortcoming identified with the current SR method - the smearing of structures on small scales - is not an issue for predictions of, e.g. neutral atomic hydrogen (HI) 21-cm intensity mapping (e.g. Villaescusa-Navarro et al., 2018) or the growth of the ionizing background during cosmological reionization (e.g. Greig et al., 2022), where the HI density field is effectively smoothed on scales of $\sim$10 Mpc. Having the means to rapidly generate matter density fields - and by extension galaxy density fields informed by high resolution hydrodyamical simulations - will be an invaluable tool for cosmology with the SKA telescope (e.g. Maartens et al., 2015).
@@ 5 Code and Data Access
All neural network models and scripts used to produce the results in this paper were implemented in Python and PyTorch and are available at https://github.com/brenjohn/DMSR-WGAN on the ”on-wgan-super-resolution” branch. Parameter files used for generating initial conditions and running the dark matter only simulations used in this work are also available there. Datasets and trained neural network models are available upon request from the corresponding author.
@@ Appendix A Appendix A: Results from Generalized Critic Models
Here we include some additional results from WGANs with the generalized critic architecture described in section 2.2.2. Three WGANs were trained whose critic models had 0, 1 and 2 additional residual layers in their density branches. The WGANs were therefore trained using density fields of 25 $\rm{h^{-1}}$ Mpc cubic patches with $32^{3}$, $68^{3}$ and $140^{3}$ cells respectively. Loss curves for the WGANs are shown in Fig. 7 where the loss function is given by equation 1. A comparison of outputs from the three models is shown in Fig. 8.
@@
Figure 8: Here we show a comparison of the same 25 $\rm{h^{-1}}$ Mpc region shown in Fig. 3, enhanced by three different DMSR-WGAN models using different resolution levels to train the critic model. From left to right, the models used had 0, 1 and 2 additional layers in the density branch of the critic model as described in the methodology. As can be seen, the presence of smeared features persists even when the resolution of the density field has been increased by 2 levels. Although the smeared features become sharper, indicating that the effect is being reduced, training WGAN models with higher density resolution quickly becomes intractable. Takeaway: Simply increasing the resolution of the density field used to train the WGAN is not a viable approach to address smearing in enhanced snapshots.

### v2

@@ 3.1 Enhanced Dark-Matter Snapshot
Figure 5: The top panel shows the halo mass function for LR, HR and SR snapshots. The halo mass function was computed using a friends-of-friends algorithm, with a linking length of $0.2$ times the mean inter-particle spacing, to identify halos. The bottom panel shows the ratio with respect to the HR halo mass function. Takeaway: The halo mass function of the enhanced snapshot closely matches that of the HR simulation, including at mass scales not resolved by the LR snapshot.
@@
The halo mass function for each snapshot was also computed using a friends-of-friends method (e.g., Davis et al., 1985) with a linking length of $0.2$ times the mean inter-particle spacing and by only considering halos with at least 100 particles. We similarly find that the enhanced data agrees with the HR data. The mass functions are shown in Fig. 5 and demonstrate that the enhanced snapshot successfully reproduces the abundance of halos across the same mass range as the HR data. In particular, the enhanced snapshot captures the higher abundance of low-mass halos, which are poorly resolved in the LR data, while maintaining accuracy for high-mass halos. These findings suggest that the SR displacement fields retain important statistical features of the target HR simulation.
@@
Enhancing LR snapshots with a trained generator model takes about 2 seconds on a single CPU  [^2] The CPU used was Intel Core i9-10940X CPU. This is negligible compared to the wall-clock time for the LR simulation to reach redshift 00 using the same CPU (approximately 7 minutes). While the wall-clock time for the HR counter part to reach the same redshift was 1 hour, generating HR snapshots with SR is significantly faster, with greater speed-ups expected when larger enhancement factors are used.
3.2 Morphology of cosmic structure
One major discrepancy we found between real and generated HR data is in relation to the shape of dark-matter halos. We found that halos in the data created by a trained generator tend to have smeared features compared to halos formed in real HR simulations. We find this discrepancy is most apparent in scatter plots of the particle positions. In Fig. 6 we show scatter plots of the same snapshots used to create the density fields in Fig. 3. The smearing of halo shapes appears to occur on length scales below the spatial resolution of the density fields used during training. As a result, the smeared features are not resolved by the cloud-in-cells density fields used during training, meaning the neural network models may learn to ignore it. This discrepancy may limit the applicability of SR methods in analyses that are sensitive to small-scale halo morphology.
To quantify this discrepancy, we computed the shape of each halo using the method denoted $E1$ in (Zemp et al., 2011). This method fits an ellipsoid to a given dark matter halo. We denote the lengths of the principle semi-axes of an ellipsoid by $a>b>c$ and the ratios $b/a$ and $c/a$ characterize the ellipsoid shape. Specifically, spherical halos have $(b/a,c/a)\approx(1,1)$, elongated (or prolate) halos have $(b/a,c/a)\approx(0,0)$ while flattened (or oblate) halos have $(b/a,c/a)\approx(1,0)$. In Fig. 7 we plot the distribution of $(b/a,c/a)$ for halos from both a real HR snapshot and a generated SR snapshot. The distributions were computed using the Gaussian kernel density estimation method from the scipy Python library. We find the SR halo shape distribution is shifted toward lower values of $(b/a,c/a)$, indicating a tendency toward more prolate structures compared to HR halos. To compute the statistical significance of this difference, we performed a permutation test using the $L^{2}$ distance between the distributions as the test statistic. The observed $L^{2}$ difference was 0.0044, which was significant at the $\alpha=0.5\%$ level, with a $p$-value of 0.2%.
Figure 7: Here we compare the distribution of shape parameters $(b/a,c/a)$ for both HR and SR halos. Contour plots of the two joint distributions are shown in the main panel while the marginal distributions are shown in the top and right hand panels. We find the distributions for SR halos are shifted towards lower values of $b/a$ and $c/a$ in comparison to the HR distributions. Takeaway: SR halos are more elongated (or prolate) compared to HR halos.
This bias towards overly smooth dark-matter distributions also affects the shape of the large-scale cosmic web. As seen in Fig. 6, the smearing effect appears to erase much of the filamentary substructure, replacing it with a more diffuse distribution of dark matter along the filaments. To quantify this morphological difference between real and generated data, we analyzed variations of the minimal spanning tree (MST) (Barrow, Bhavsar & Sonoda, 1985; Naidoo et al., 2020) statistic. MSTs were constructed from the dark-matter particle positions in both HR and SR snapshots using the Kruskal algorithm as implemented in scipy. The MSTs were also pruned of all branches with 25 or fewer links. Any edges longer than 3 times the average edge length were also removed. This level of pruning was enough to remove almost all of the branches extending along the filaments of the HR data, leaving dense clusters of MST nodes corresponding to the nodes of the cosmic web. In contrast, large branches corresponding to filaments persist in the SR data after applying the same pruning criteria. Projections of the pruned MSTs for both cases are shown in Fig. 8. We also present the edge length distributions of the MSTs in the top panel of Fig. 9 and the distribution of the number of edges in a branch in the lower panel. Compared to HR data, SR data exhibits both longer edge lengths and branches composed of more edges. Both of these are indicative of a more diffuse distribution of dark-matter particles along the cosmic web, particularly along the filaments. These results reinforce the conclusion that the SR method struggles to accurately reproduce the geometry of the dark-matter distribution.
Figure 8: Projections of the MSTs for the HR and SR data are shown on the left and right respectively. After pruning and separating, the HR MST consists of 521 connected components, all of which represent nodes of the cosmic web. In contrast, the SR MST has 1017 connected components that exhibit prominent branches that still extend along the filaments. Takeaway: The SR MST retains extended branches even after applying a pruning criterion that removes them from the HR MST.
Figure 9: The distribution of edge lengths for both HR and SR MSTs are shown in the top panel. The bottom panel shows the distribution of the number of edges in a branch for both MSTs. Takeaway: In comparison to the HR MST, the SR MST exhibits longer edge lengths and branches consisting of a greater number of edges.
@@ 3.2 Morphology of cosmic structure
This smearing effect is consistent with known limitations in the SR literature where some GAN-based models can produce outputs that are biased toward overly smooth solutions (Wang et al., 2023, e.g.,). While in some cases this bias can be attributed to certain loss functions (Blau & Michaeli, 2018; Wang et al., 2019, e.g.,), a key issue here is that the output of our WGAN model is inherently under-constrained on scales below the resolution of the training targets (such as the cloud-in-cells density field). To empirically test whether increasing the spatial resolution of the discriminator input could mitigate this effect, we trained two additional models using the generalized critic architecture described in Section 2.2.2. These models used density fields with spatial dimensions of $68^{3}$ and $140^{3}$, corresponding to one and two additional residual blocks in the density branch, respectively. The base number of channels for the density branch in both cases was set to $16$ while it was kept at $64$ for the main branch. A comparison of dark-matter distributions generated by these models is shown in Fig 11 in the appendix. We found that increasing the resolution of the density field in this way did not remove the smearing effect. Moreover, the wall-clock time per training epoch increased by factors of 2 and 140, respectively. Further increases in resolution were not tested, as the required memory exceeded the capacity of the H100 GPU used for this work.
@@ 4 Conclusions and Outlook
In this work we trained several WGAN models to enhance the particle resolution of snapshots from dark-matter-only simulations, following the method outlined in Li et al. (2021). While we found that these models could produce data with summary statistics that agree well with real high-resolution (HR) simulations, we identified a notable limitation: the shapes of dark matter halos in the super-resolved (SR) data are biased by smeared features not present in the real HR data. This discrepancy in halo morphology, visible in both scatter plots and halo shape metrics, suggests that while global statistics may be preserved, important small-scale structural details are lost. This limitation must be considered when using SR models for scientific analyses that are sensitive to halo shapes.
@@
Despite these limitations, SR methods remain a valuable tool for cosmological studies. Given the impressive demonstrations of the techniques used in Zhang et al. (2024) and Jacobus et al. (2024), it is tempting to judge these methods as cheap replacements for HR simulations. Nevertheless, it is important to recognize that these models are approximating the physics occurring on length scales that are not resolved by the LR simulation. Therefore, rather than being viewed as a replacement for HR simulations, SR methods should be regarded as simulation-informed subgrid models. As such, the value of these methods stems from their ability to help bridge the gap between LR and HR simulations when computational resources are limited.
@@
Future work could address the issue of smeared features by developing additional loss terms that account for physical properties beyond summary statistics. For example, quantifying the divergence of halo shapes from those appearing in HR data could provide a regularization term to discourage the WGAN from generating overly smoothed solutions. To this end, the shape parameters used in this work or other metrics, such as Minkowski functionals, that encode information about dark matter halos could be useful. This approach aligns with the methods used by Helfer et al. (2024) to correct numerical relativity simulations that break conservation laws during mesh refinement. Exploring alternative generative frameworks, such as stable diffusion models, may also provide insights into whether similar limitations persist or if new architectures can overcome this issue. A detailed comparison between WGANs and stable diffusion methods in this context could help identify the strengths and weaknesses of each approach.
@@
Applications of SR techniques are significant, and we will explore some of those applications in future works. One example would be the ability to extend the resolution of cosmological N-body simulations to smaller scales opens the possibility of better tracking the assembly history of halos at a relatively small cost. In addition, to the positional information used in this work, this would require temporal (redshift) information and kinematics - however both of these extensions are tractable. Gaining insight into the assembly history of halos at high redshift would be particularly useful for understanding the formation of (rare) early cosmic structures which require both large volume statistics and high resolution (e.g., Trinca et al., 2022). These smaller halos (atomic cooling halos and molecular cooled halos) have the potential to host the first stellar objects in our Universe. Semi-analytical models, which use merger trees as their backbone, could benefit hugely from SR techniques (e.g., Spinoso et al., 2023).
@@
A second example is the use of SR techniques to capture small-scale structure within simulations of large-scale structure spanning several gigaparsecs. The shortcoming identified with the current SR method may not be problematic for certain applications in this direction. For instance, neutral atomic hydrogen (HI) 21-cm intensity mapping (e.g., Villaescusa-Navarro et al., 2018) or modeling the growth of the ionizing background during reionization (e.g., Greig et al., 2022), since the HI density field is effectively smoothed on scales of $\sim$10 Mpc. The ability to rapidly generate matter density fields - and, by extension, galaxy density fields informed by high-resolution hydrodynamical simulations - will be an invaluable tool for cosmology with the SKA telescope (e.g., Maartens et al., 2015).
@@ 5 Code and Data Access
All neural network models and scripts used to produce the results in this paper were implemented in Python and PyTorch (Paszke et al., 2019) and are available at https://github.com/brenjohn/DMSR-WGAN on the ”on-wgan-super-resolution” branch. Steps for reproducing results are also available there along with model weights used for this work. For recreating the datasets, the parameter files used for generating the initial conditions and running the dark-matter-only simulations, from which the datasets were created, are also provided. The work presented here made use of the numpy, scipy, networkx and matplotlib Python libraries (Harris et al., 2020; Virtanen et al., 2020; Hagberg, Schult & Swart, 2008; Hunter, 2007).
@@ Appendix A Appendix A: Results from Generalized Critic Models
Here we include some additional results from WGANs with the generalized critic architecture described in section 2.2.2. Three WGANs were trained whose critic models had 0, 1 and 2 additional residual layers in their density branches. The WGANs were therefore trained using density fields of 25 $\rm{h^{-1}}$ Mpc cubic patches with $32^{3}$, $68^{3}$ and $140^{3}$ cells respectively. Loss curves for the WGANs are shown in Fig. 10 where the loss function is given by equation 1. A comparison of outputs from the three models is shown in Fig. 11. Note that in Wasserstein GANs, the absolute value of the critic loss approximates the Wasserstein distance between the generator’s learned distribution and the true distribution of the training data. Therefore, lower (i.e., closer to zero) critic loss values generally indicate a smaller discrepancy between the two distributions. For a detailed discussion, we refer the reader to Gulrajani et al. (2017). However, it is important to emphasize that the critic loss is not typically used as a direct measure of generator quality. Consequently, loss values are not used for model selection or early stopping in WGAN training. Instead, critic and generator loss curves serve as diagnostic tools to monitor training dynamics and assess stability. We include these loss curves in the appendix to assist readers who may wish to reproduce our results or compare training behavior.
@@
Figure 11: Here we show a comparison of the same 25 $\rm{h^{-1}}$ Mpc region shown in Fig. 3, enhanced by three different WGAN models using different resolution levels to train the critic model. From left to right, the models used had 0, 1 and 2 additional layers in the density branch of the critic model as described in the methodology. As can be seen, the presence of smeared features persists even when the resolution of the density field has been increased by 2 levels. Although the smeared features become sharper, indicating that the effect is being reduced, training WGAN models with higher density resolution quickly becomes intractable. Takeaway: Simply increasing the resolution of the density field used to train the WGAN is not a viable approach to address smearing in enhanced snapshots.